OpenStudy (anonymous):

f(x)=Sqrt{x}/(Sqrt{x}-3),find the inverse and domain and range of this function.

OpenStudy (anonymous):

\[f(x)=\frac{\sqrt{x}}{\sqrt{x}-3}\] like this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

domain. \[x\geq 0,x\neq 9\] clear yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

range i believe is all real numbers except of course 1

OpenStudy (anonymous):

for inverse put \[x=\frac{\sqrt{y}}{\sqrt{y}-3}\] and solve for y

OpenStudy (anonymous):

thats where my problem is i am having problem solving for or getting the inverse

OpenStudy (anonymous):

\[(\sqrt{y}-3)x=\sqrt{y}\] \[x\sqrt{y}-3x=\sqrt{y}\] \[x\sqrt{y}-\sqrt{y}=3x\] \[\sqrt{y}(x-1)=3x\] \[\sqrt{y}=\frac{3x}{x-1}\]

OpenStudy (anonymous):

Then square both sides right?

OpenStudy (anonymous):

then square. check my algebra because it is late

OpenStudy (anonymous):

how is range all real numbers expect 1...can u pls explain it...

OpenStudy (anonymous):

looks good

OpenStudy (anonymous):

i get \[f^{-1}(x)=\frac{9x^2}{(x-1)^2}\]

OpenStudy (anonymous):

well i think it may be a calc problem but one thing is for sure. a fraction is only when when the numerator and the denominator are the same. and that is impossible here, by inspection

OpenStudy (anonymous):

i meant to say a fraction is only ONE when the top and bottom are equal

OpenStudy (anonymous):

this thing has a vertical asymptote at 3, and it goes to minus infinity from one direction and positive infinity from the other. look at a picture and you will see it. but it can never be 1

OpenStudy (anonymous):

i got vertical asymptote at 9...and picture on my calculator is at 4th quadrant only...how to get correct picture

myininaya (myininaya):

what satellite meant to say is there is a vertical asymptote at x=9 not 3 he made a type-0

OpenStudy (anonymous):

but how is range all real numbers except 1... i still cant get that..

OpenStudy (anonymous):

i dont get it either

OpenStudy (anonymous):

i dont get it either

OpenStudy (anonymous):

just believe i am sure it is right

OpenStudy (anonymous):

now you are calling them typos. see!

OpenStudy (anonymous):

ok first of all it is clear that f cannot be 1. now you have a formula for the inverse. so what else can't f be?

OpenStudy (anonymous):

the domain of the inverse is all real numbers except 1. that is because f cannot be 1!

OpenStudy (anonymous):

but f can be anything else you like and your inverse proves it. suppose i want f to be 7. ok i just take \[f^{-1}(7)\] and that finds the x that gives it to me

OpenStudy (anonymous):

so i say the range of f is all numbers except 1, and if you give me any other number z i will take \[f^{-1}(z)\] and say look that is the x that makes \[f(x)=z\]

OpenStudy (anonymous):

ohhk i got it...btw is there any way we can determine range jus by looking at the function...not its inverse..

OpenStudy (anonymous):

i have seen people do it by "solving for x" which is basically finding the inverse. calc is easier. but you might find that the inverse involves a square root and then you have to make sure what is inside is positive. myininaya can help you with that. i just graph

OpenStudy (anonymous):

so the range would be all numbers grater than or equal to 0 but not 9

OpenStudy (anonymous):

?

OpenStudy (anonymous):

hey thanks......btw can we see ur graph...

OpenStudy (anonymous):

what is the range, is it the domain of the orignal?

OpenStudy (anonymous):

what sis the range?

OpenStudy (anonymous):

WHAT SI THE rNAGE?

OpenStudy (anonymous):

so range is all numbers except 1 for the original function,also can be said as domain of the inverse

OpenStudy (anonymous):

but what is range of inverse?

OpenStudy (anonymous):

range of inverse is domain of the original function

OpenStudy (anonymous):

all real numbers except 9

OpenStudy (anonymous):

hey man do you get related reates

OpenStudy (anonymous):

still on quiz qs....

OpenStudy (anonymous):

but it cant be all reall numbers, it has to numbers greater than o equal to zero except 9

OpenStudy (anonymous):

havent got that far...

OpenStudy (anonymous):

i am done with it

OpenStudy (anonymous):

look at this post and click on it, i will give them to you

OpenStudy (anonymous):

http://www.twiddla.com/582353